OpenStudy (anonymous):

help please Currently, Sam's coach is three times older than Sam. When Sam started soccer 15 years ago, his coach was six times his age. How old is Sam now? 15 10 25 20

hartnn (hartnn):

Let Sam's age now be x Sam's coach is three times older than Sam, so, his coach's age now = 3x 15 years ago, Sam's age = x-15 his coach's age = 3x -15 , then his coach was six times his age, so 6(x-15) = 3x-15 can you solve this equation to get the value of x ?

OpenStudy (anonymous):

@hartnn -10?

hartnn (hartnn):

how can age be negative ? :P

hartnn (hartnn):

or did you mean age = 10 ?

hartnn (hartnn):

anyways 10 is incorrect.

hartnn (hartnn):

did you try solving this ? 6(x-15) = 3x-15